OpenStudy (anonymous):

how about this one 20% writen as a fraction

OpenStudy (anonymous):

"per centum" means "per 100." So 20 percent would be 20 per 100, or \[\frac{20}{100}\] This can be reduced to 1/5

OpenStudy (lalaly):

20/100 = 1/5